given that a root is two times the other root of the quadratic equation 4x^2-18x=k=0.
find the value of k

Answer by ankor@dixie-net.com(22740)

given that a root is two times the other root of the quadratic equation
4x^2 - 18x + k = 0.
find the value of k
:
Let a = one root
then
2a = the other root
:
A couple factors could be
(2x-a)(2x-2a) = 0
then using the middle term coefficient
2a + 2(2a) = -18
2a + 4a = -18
6a = -18
a = -3
then
2a = -6
the factors
(2x-3)(2x-6) = 0
FOIL
4x^2 - 12x - 6x + 18 = 0
4x^2 - 18x + 18 = 0
k = 18